OpenStudy (anonymous):

2log base 4x-log base 4(x+5)=2

OpenStudy (zzr0ck3r):

\[2log_4(x)-log_4(x+5) = 2?\]

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

\[2log_4(x)-log_4(x+5) = log_4(x^2) - log_4(x+5) = \]

OpenStudy (zzr0ck3r):

do you get this step?

OpenStudy (anonymous):

Yes thx

OpenStudy (zzr0ck3r):

\[log(a)-log(b) = log(\frac{a}{b})\]

OpenStudy (zzr0ck3r):

so \[log(x^2) - log(x+5)= log(\frac{x^2}{x+5})\] so \[log_4(\frac{x^2}{x+5})=2\] this means \[4^2 = \frac{x^2}{x+5}\] so \[16(x+5) = x^2 \rightarrow x^2-16x-80=0\] can you solve \[x^2-16x-80=0?\]

OpenStudy (zzr0ck3r):

(x-20)(x+4)=0 x=20 and x=-4 but \[log_4(x) \] is not defined for x<= 0 so, we take x=20 to be the solution

OpenStudy (anonymous):

Thanks!